OpenStudy (anonymous):

Which class of numbers is described by the set given below? {0, 1, 2, 3, ... } A. rational numbers B. irrational numbers C. whole numbers D. integers

OpenStudy (zzr0ck3r):

I would Google the definition of each.

OpenStudy (zzr0ck3r):

this is how I remember it the natural numbers \(\mathbb{N}\) are the counting numbers, I think "nature" ....fingers, 1,2,3,5,6,7...... then we have whole numbers, the way I remember this is "whole" looks like "wh0le" so i think 0,1,2,3,4,5,6,7,...... then there the negative numbers we need to deal with , first letter of Integer, looks like a sideways - this makes me think negative goes with integers, and we call these \(\mathbb{Z}\) now we have rationals, well those are things that can be written as "ratios" of integers a/b (where b is not 0), we call these \(\mathbb{Q}\) then we have irational numbers, i.e. the things that cant be written as a fraction of integers (like \(\sqrt{2}, \text{ or } \pi\)) if we put all these together we get the real numbers \(\mathbb{R}\).

OpenStudy (zzr0ck3r):

^^ travel into the mind of the dyslexic mathematician to see the silly ways one might remember things:)

OpenStudy (zzr0ck3r):

I should have said the negatives and the wholes together give the integers \(\mathbb{Z}={.....-3,-2,-1,0,1,2,3,.......}\)

OpenStudy (zzr0ck3r):

So what is the answer to your question?